OpenStudy (anonymous):

How do the graph of y=√4x and y=2√x compare?

OpenStudy (anonymous):

y=√x to y=√4x is a horizontal compression by a factor of 1/4. y=√x to y=2√x is a vertical expansion by a factor of 2. How are these two graphs equal?

OpenStudy (campbell_st):

go and download geogebra... a free graphing geometry package... they are one in the same line \[y = \sqrt{4 \times x} = \sqrt{4}\times \sqrt{x} = 2\sqrt{x}\] just show the congruency via the algebra But I have attached a copy of the graphs